How many grams of CO2 are produced when 100 g C4H10 undergoes combustion with 200 g O2?


Use the following molar masses: C4H10=58g/mol, O2=32 g/mol and CO2=44g/mol.

Compare the moles of product formed from each reactant.


1
Expert's answer
2022-03-16T15:39:03-0400

M(C4H10) = 58 g/mol;

M(O2) = 32 g/mol;

M(CO2) = 44/mol;

M(H2O) = 18 g/mol;

m(C4H10) = 100 g;

m(O2) = 200 g;

2C4H10 + 13O2 = 8CO2 + 10H2O;

n(C4H10) = m(C4H10)/M(C4H10) = 100/58 = 1.72 mol;

n(O2) = m(O2)/M(O2) = 200/32 = 6.25 mol.

Given the stoichiometric coefficients:

n(C4H10)' = n(C4H10)/2 = 0.86 mol;

n(O2)' = n(O2)/13 = 6.25/13 = 0.48 mol;

n(C4H10)' is higher than n(O2)', therefore C4H10 is excess and O2 is limited reagent.

For chemical reaction:

n(H2O) = 10/13 * n(O2) = 10 * 6.25/13 = 4.81 mol;

n(CO2) = 8/13 * n(O2) = 8 * 6.25/13 = 3.85 mol;

m(CO2) = n(CO2) * M(CO2) = 3.85 * 44 = 169.4 g.

Answer: 169.4 g.